OpenStudy (anonymous):

The line of symmetry for the quadratic equation y=ax²-8x-3 is x=2. What is the value of ¨a¨? A. -2 B. -1 C. 2 I had a teacher try and do this and they couldnt do it. s help me please!!!!

OpenStudy (anonymous):

go to mathway.com

OpenStudy (er.mohd.amir):

Q is wrong may be this is right y=ax^2-8x+8 at x=2 y=0 then 4a-16+8=0 4x=8 x=2

OpenStudy (er.mohd.amir):

or x=-b/2a here b=-8 a=a x=2 2=-(-8)/2a 4a=8 a=2

OpenStudy (bee_see):